# The Expressive Capacity of State Space Models: A Formal Language Perspective

### v2

@@ 5 Discussion
Our work belongs to an incipient line of research into the expressiveness of SSMs [33, 49]. It is closely related to a long string of work studying the expressive capacity of neural sequence models, which has so far focused on recurrent networks [e.g. 62, 6, 28] and, more recently, self attention [e.g. 10, 48, 66]. A second link is to the classical and long-standing study of linear dynamical systems and control theory [34]. For instance, Theorem 2 relies the asymptotic convergence of an SSM on certain inputs, establishing a link to the asymptotics of linear systems [e.g. 55].
@@
While theoretical in nature, our results have several actionable implications for SSM and LLM research, informing the rapidly growing research on SSM-based LLMs. First, encouragingly, SSMs can keep track of bounded hierarchical structure with optimal memory even without explicitly implementing a stack (Theorem 6), suggesting that simple diagonal linear state updates may be sufficiently powerful for modeling the hierarchical structure of language. Second, SSMs resolve a basic failure mode of self-attention in flip-flop state tracking while being parallellizable (Theorem 1). Overall, SSMs and attention have overlapping but distinct strengths. This lends support to the development of hybrid architectures interleaving SSM and attention layers, as instantiated very recently by Jamba [42]. Third, nonnegative gates as obtained by exponential or sigmoid parameterizations provably restrict expressive capacity, even in non-time-invariant SSMs (Theorem 2). While [23] found no evidence that complex-valued paramerizations improved over real-valued ones in the language modality, our results suggest revisiting this question, at least for tasks where periodic state-tracking abilities may be important. Fourth, while exactly characterizing the capacity of transformers has proven difficult even in the finite-state case, Theorem 4 provides a decidable characterization of the regular languages – equivalently, finite-state tracking problems – that SSMs such as Mamba can model. Such decidable characterizations may make it easier to theoretically predict abilities and anticipate failures of LLMs; exploring the implications of this characterization in more realistic setups is an exciting direction for future research.
@@ Acknowledgments
We thank Mark Rofin for useful discussion about Theorem 2.
@@ Appendix A Instantiations of General Framework in SSM Models
Here, we survey how (1) is instantiated in a range of SSMs. As stated in Section 2, we refer to SSMs where the gate $A$ does not depend on $x_{t}$ as time-invariant. An equivalent terminology is the distinction between “Weak Linear Time Invariant Convolutional Models” (i.e., time-invariant) and “Linear Time Variant Models” (i.e., non-time-invariant) in Akyürek et al. [1].
@@ A.1 Non-Time-Invariant Models
Approximately simultaneously with or more recently than [23], a range of non-time-invariant SSMs have been introduced [14, 72, 57, 56]. This category also covers highly similar earlier RNN variants [8, 41].
@@
In Mamba, (2) and (3) directly map onto Eqs. (2a) and (2b) in Gu and Dao [23]. The notation of Gu and Dao [23] use a matrix multiplication $\overline{A}h_{t-1}$ instead of elementwise multiplication $A(x_{t})\circ h_{t-1}$ in (REF), but importantly, Mamba’s $\overline{A}$ is diagonal, so we can take $A(x_{t})_{i}=\overline{A}_{ii}$. Due to exponential parameterization, its entries are nonnegative.
@@
The RG-LRU layer of Griffin [14] uses the equation
@@
where $a_{t},i_{t}$ are neurally parameterized in terms of $x_{t}$ but not $h_{<t}$; by design, $a_{t}\in(0,1)$. $\phi$ is instantiated in terms of linear transformations, GeLU, and RMSNorm (Figure 2 in De et al. [14]). The local attention used by Griffin can be subsumed into an SSM layer (Remark 18).
Gated Linear Attention [GLA 72]
This model (Section 4.4 in Yang et al. [72]) instantiates our framework using a recurrence of the form (1); while the state is two-dimensional in this model, the update is performed by elementwise products as in (1). The gate is obtained by applying sigmoid to a linear transformation of $x_{t}$; thus, its entries are in $(0,1)$. $\phi$ is instantiated in terms of SwiGLU and LayerNorm.
@@
HGRN [57] and HGRN2 [56] are defined by a recurrence of the form (1); the gate entries are $\in(0,1)$ by design. $\phi$ is instantiated in terms of GLU, linear transformations, and normalization. In HGRN, the state is complex, but crucially the gate remains real-valued.
@@ A.2 Time-Invariant Models
Time-invariant SSMs introduced before late 2023 are surveyed by Gu and Dao [23, Appendix B], such as [47, 67, 52]. Time-invariant SSMs have often used complex-valued states and gates; this does not have a major impact on our results: First, as complex-valued SSMs subsume real-valued ones, our positive results carry over. Second, our negative result about PARITY is affected by this distinction and requires a separate argument, see Theorem 13.
Note also that $Ah_{t-1}$ is often described as a general matrix multiplication, but $A$ is diagonalizable (e.g. Lemma 3.2 in Gu et al. [24]; [67] for RetNet), which —even though implementation may be based on non-diagonalized representations [24]—renders the model equivalent to one where $A$ is diagonal from the start. This equivalence is shown as Lemma 3.1 in Gu et al. [24].
@@ B.1 Flip Flop
We begin by introducing key notions of automata theory. References for automata theory include Eilenberg [16], Hopcroft et al. [30], Sakarovitch [58]. We will provide those key notions that are necessary to prove our results. We will focus on deterministic finite-state-automata (DFA), and simply refer to them as finite-state-automata. [^4] A closely related notion is the semiautomaton, which is the notion considered in the closely related work Liu et al. [44]. Semiautomata lack a fixed start state $q_{0}$. We include $q_{0}$, but this difference is not substantial for our formal results. First,
@@
Kleene’s Theorem [37] asserts that a language $L\subseteq\Sigma^{*}$ is regular (i.e., defined by a regular expression) if and only if it is recognized by some finite-state automaton.
@@
A set-reset automaton is a finite-state-automaton where $(Q\textbackslash\{q_{0}\})\subseteq\Sigma$ and
@@
Intuitively, such an automaton keeps recording the last seen symbol from a designed set $Q\subseteq\Sigma$. Such an automaton is easily simulated with a single non-time-invariant SSM layer:
@@
Theorem 10.
(Repeated from Theorem 1) A two-layer SSM with finite precision can predictively model $\mathcal{L}_{FF}$ at arbitrary lengths.
@@ B.2 Difficulty of Representing PARITY
Theorem 12.
(Repeated from Theorem 2) No SSM satisfying nonnegative can recognize PARITY at arbitrary input lengths with finite precision.
@@
We now need to understand the behavior of $\operatorname{Mix}_{1}(h_{t},x_{t})$. Recall that, based on our survey (Appendix A), we allowed it to contain linear, GLU [13], and SwiGLU [61] components. If $\operatorname{Mix}_{1}(h_{t},x_{t})$ implements a linear transformation only, each entry likewise may converge, diverge linearly, or diverge exponentially. We note that—if $\sigma$ is the sigmoid function—$\sigma(u_{t})$ always converges, as $\sigma$ simply saturates to 0 or 1 if $u_{t}$ diverges. Hence, if $\operatorname{Mix}_{1}(h_{t},x_{t})$ implements GLU, each entry likewise may converge, diverge linearly, or diverge exponentially. Finally, if $\operatorname{Mix}_{1}(h_{t},x_{t})$ implements SwiGLU, each entry of the result will be a product of a linear combination of the form $u_{t}$, and $Swish_{\beta}$ applied to another such linear combination. Depending on the behavior of these two $u_{t}$-like terms, the outcome will behave as a product of sequences that may converge exponentially, diverge exponentially, or diverge linearly – e.g., the outcome may also diverge quadratically, or converge as $n\alpha^{-n}$, etc.
[…]
@@
For any non-star-free regular language $\mathcal{L}$, there are words $u,v,w$ such that the membership $uv^{n}w\in\mathcal{L}$ is determined by the parity of $n$ [46]. Fix any such $u,v,w\in\Sigma^{*}$.
@@
We now simply absorb the operation $X\mapsto\psi_{|w|}\circ X+\gamma_{|w|}$ into $\operatorname{Mix}_{2}$, and obtain by the same arguments as in the proof of Theorem 2 that $z_{t|v|+|w|}^{(L)}$ converges as $r\rightarrow\infty$. This is a contradiction to the claim that the parity of $t$ can be read out from $z_{t|v|+|w|}^{(L)}$ at finite precision. ∎
@@ B.3 Proof of Theorem 4
Our proof of Theorem 4 will rely on the algebraic theory of finite automata, specifically the cascade product and the Krohn-Rhodes Theorem [39]. These techniques, originally developed in the 1960s, have recently been introduced to the theoretical study of transformers by Liu et al. [44]; we provide self-contained definitions and somewhat different notation, tailored to our proofs about state-space models. In general, we will find that the properties of state-space models allow more natural and directly length-generalizing implementations of these algebraic notions than what is possible for transfomers.
Recall the definition of a finite-state-automaton (Definition 7). Our construction will build on an important operation on automata, the cascade product [39, 16, 22]:
@@
While the formal definition is cumbersome, the intuition behind it is simple: The cascade product corresponds to first reading a word ${\bf w}$ with $\mathcal{A}_{1}$, recording the state sequence $q_{0},q_{1},\dots,q_{|{\bf w}|}\in Q_{1}$ and – at each $t=1,\dots,|{\bf w}|$ – pasting the state $q_{t-1}$ together with the input symbol $w_{t}\in\Sigma_{1}$ – resulting in a word over a new alphabet $Q_{1}\times\Sigma_{1}$, and then running $A_{2}$ on the resulting word. The overall state of $A_{2}\wr A_{1}$ after reading a word is the tuple of the states reached by $A_{2}$ and $A_{1}$. Note that we write $A_{2}\wr A_{1}$, rather than, $A_{1}\wr A_{2}$, because the second argument of the cascade product ($A_{1}$) intuitively reads the input first, preprocessing it for the other automaton, $A_{2}$.
@@
Fact 16 (Consequence of Krohn-Rhodes Theorem [39] and Schützenberger’s Theorem [60]).
@@
We now formally show that cascade products can be translated to SSM stacking. We need an auxiliary lemma, which provides a single-layer SSM that encodes the input $w_{t-1}$ in state $h_{t}$ – we will use it to forward information about the state of $\mathcal{A}_{1}$ at $t-1$ to $\mathcal{A}_{2}$ at $t$:
@@
Now we separately consider the state $h_{t}$ depending on the form of the prefix $w_{1\dots t}$: If $w_{1\dots t}=\dots 00$, then
@@
If $w_{1\dots t}=\dots 10$, then
@@
Now, for each $\sigma\in\Sigma$, we run such a one-layer SSM where 00 represents $\sigma$ and $1$ represents all other characters. [^6] In fact, using a binary encoding of $\Sigma$, one can achieve $d=4\log|\Sigma|$. By running these in parallel, we obtain an SSM with $d=4|\Sigma|$ from whose states one can read out $w_{t-1}$ at finite precision. As the entries in $h_{t}$ are all bounded by $2$, we find $\|h_{t}\|_{2}\leq 2\sqrt{d}$ independent of $t$, and the margin is still bounded away from zero after normalization, and thus in $z_{t}$, where we can assume $\operatorname{Mix}_{1}$, $\operatorname{Mix}_{2}$ to be the identity. ∎
@@
Some SSMs include local convolutions [e.g. 21, 23] or local attention [14], which aggregate information from a local window of some width $\Delta>0$. These do not increase the expressive capacity beyond SSMs as we have defined in (1-2), as aggregation of local information can be simulated with a single SSM layer: Using the layer constructed in the proof of Lemma 17, given the state $h_{t}$, once one has read out $w_{t-1}$ as described in the proof, one can recover $h_{t-1}$ from $h_{t}$ and $x_{t}$; then inductively read out $w_{t-2}$ using $h_{t-1}$ and $x_{t-1}$, etc. Thus, up to any given width $\Delta>0$, one can read out $w_{t-\Delta},\dots,w_{t-1}$ from the state $h_{t}$ of this layer at finite precision.
@@
We note that a conceptually related result holds for transformers [Lemma 12 in 44]. However, SSMs allow a simpler and length-independent construction, as they do not require positional encodings to implement such a construction.
@@
Theorem 20 (Repeated from Theorem 4).
Assuming Nonnegative, SSMs can predictively model a regular language $\mathcal{L}$ at finite precision if and only if $\mathcal{L}$ is star-free.
@@
Assume $\mathcal{L}$ is star-free. By the Krohn-Rhodes theorem, there is an automaton $\mathcal{A}$ that is a cascade product of some set-reset automata that recognizes $\mathcal{L}$. By Lemmas 9 and 19, there is an SSM that computes the state sequence of that automaton.
@@
The set of regular languages known to be in $\operatorname{TC}^{0}$ is the set of regular languages whose syntactic monoid contains no non-solvable groups [5]. These languages are recognized by cascade products of set-reset automata and automata perfoming modular counting [64]. By the remark above, together with Lemma 9 and Lemma 19, such cascade products can be simulated by SSMs. ∎

### v3

@@ 5 Discussion
Our work belongs to an incipient line of research into the expressiveness of SSMs (jelassi2024repeat; merrill2024illusion). It is closely related to a long string of work studying the expressive capacity of neural sequence models, which has so far focused on recurrent networks (e.g. siegelman1991neural; bhattamishra2020ability; hewitt2020rnns) and, more recently, self attention (e.g. chiang2023tighter; merrill2023logic; strobl2023survey). A second link is to the classical and long-standing study of linear dynamical systems and control theory (kalman1960general). For instance, Theorem 2 relies the asymptotic convergence of an SSM on certain inputs, establishing a link to the asymptotics of linear systems (e.g. phillips1992asymptotics).
@@
While theoretical in nature, our results have several actionable implications for SSM and LLM research, informing the rapidly growing research on SSM-based LLMs. First, encouragingly, SSMs can keep track of bounded hierarchical structure with optimal memory even without explicitly implementing a stack (Theorem 6), suggesting that simple diagonal linear state updates may be sufficiently powerful for modeling the hierarchical structure of language. Second, SSMs resolve a basic failure mode of self-attention in flip-flop state tracking while being parallellizable (Theorem 1). Overall, SSMs and attention have overlapping but distinct strengths. This lends support to the development of hybrid architectures interleaving SSM and attention layers, as instantiated very recently by Jamba (lieber2024jamba). Third, nonnegative gates as obtained by exponential or sigmoid parameterizations provably restrict expressive capacity, even in non-time-invariant SSMs (Theorem 2). While Gu2023Mamba found no evidence that complex-valued paramerizations improved over real-valued ones in the language modality, our results suggest revisiting this question, at least for tasks where periodic state-tracking abilities may be important. Fourth, while exactly characterizing the capacity of transformers has proven difficult even in the finite-state case, Theorem 4 provides a decidable characterization of the regular languages – equivalently, finite-state tracking problems – that SSMs such as Mamba can model. Such decidable characterizations may make it easier to theoretically predict abilities and anticipate failures of LLMs; exploring the implications of this characterization in more realistic setups is an exciting direction for future research.
@@ Acknowledgments
We thank Mark Rofin for useful discussion about Theorem 2 and we thank anonymous reviewers for their helpful feedback. We gratefully acknowledge the insightful discussions with members of the FLaNN community which contributed to the ideas of this project. Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 232722074 – SFB 1102.
@@ Appendix A Instantiations of General Framework in SSM Models
Here, we survey how (1) is instantiated in a range of SSMs. As stated in Section 2.1, we refer to SSMs where the gate $A$ does not depend on $x_{t}$ as time-invariant. An equivalent terminology is the distinction between “Weak Linear Time Invariant Convolutional Models” (i.e., time-invariant) and “Linear Time Variant Models” (i.e., non-time-invariant) in akyurek2024context.
@@ A.1 Non-Time-Invariant Models
Approximately simultaneously with or more recently than Gu2023Mamba, a range of non-time-invariant SSMs have been introduced [De2024Griffin, yang2023gated, qin2024hierarchically, qin2024hgrn2]. This category also covers highly similar earlier RNN variants [bradbury2016quasi, lei2017simple].
@@
In Mamba, (2) and (3) directly map onto Eqs. (2a) and (2b) in Gu2023Mamba. The notation of Gu2023Mamba use a matrix multiplication $\overline{A}h_{t-1}$ instead of elementwise multiplication $A(x_{t})\circ h_{t-1}$ in (REF), but importantly, Mamba’s $\overline{A}$ is diagonal, so we can take $A(x_{t})_{i}=\overline{A}_{ii}$. Due to exponential parameterization, its entries are nonnegative.
@@
The RG-LRU layer of Griffin [De2024Griffin] uses the equation
@@
where $a_{t},i_{t}$ are neurally parameterized in terms of $x_{t}$ but not $h_{<t}$; by design, $a_{t}\in(0,1)$. $\phi$ is instantiated in terms of linear transformations, GeLU, and RMSNorm (Figure 2 in De2024Griffin). The local attention used by Griffin can be subsumed into an SSM layer (Remark 19).
Gated Linear Attention [GLA yang2023gated]
This model (Section 4.4 in yang2023gated) instantiates our framework using a recurrence of the form (1); while the state is two-dimensional in this model, the update is performed by elementwise products as in (1). The gate is obtained by applying sigmoid to a linear transformation of $x_{t}$; thus, its entries are in $(0,1)$. $\phi$ is instantiated in terms of SwiGLU and LayerNorm.
@@
HGRN [qin2024hierarchically] and HGRN2 [qin2024hgrn2] are defined by a recurrence of the form (1); the gate entries are $\in(0,1)$ by design. $\phi$ is instantiated in terms of GLU, linear transformations, and normalization. In HGRN, the state is complex, but crucially the gate remains real-valued.
@@ A.2 Time-Invariant Models
Time-invariant SSMs introduced before late 2023 are surveyed by Gu2023Mamba, such as [mehta2022long, sun2023retentive, orvieto2023resurrecting]. Time-invariant SSMs have often used complex-valued states and gates; this does not have a major impact on our results: First, as complex-valued SSMs subsume real-valued ones, our positive results carry over. Second, our negative result about PARITY is affected by this distinction and requires a separate argument, see Theorem 13.
Note also that $Ah_{t-1}$ is often described as a general matrix multiplication, but $A$ is diagonalizable (e.g. Lemma 3.2 in gu2021efficiently; sun2023retentive for RetNet), which —even though implementation may be based on non-diagonalized representations [gu2021efficiently]—renders the model equivalent to one where $A$ is diagonal from the start. This equivalence is shown as Lemma 3.1 in gu2021efficiently.
@@ B.1 Flip Flop
We begin by introducing key notions of automata theory. References for automata theory include eilenberg1974automata, hopcroft2001introduction, sakarovitch2009elements. We will provide those key notions that are necessary to prove our results. We will focus on deterministic finite-state-automata (DFA), and simply refer to them as finite-state-automata. [^5] A closely related notion is the semiautomaton, which is the notion considered in the closely related work liu2022transformers. Semiautomata lack a fixed start state $q_{0}$. We include $q_{0}$, but this difference is not substantial for our formal results. First,
@@
Kleene’s Theorem [kleene1951representationof] asserts that a language $L\subseteq\Sigma^{*}$ is regular (i.e., defined by a regular expression) if and only if it is recognized by some finite-state automaton.
@@
[…]
@@
Intuitively, such an automaton keeps recording the last seen symbol from a designated set $Q\subseteq\Sigma$. Such an automaton is easily simulated with a single non-time-invariant SSM layer:
@@
Theorem 10 (Restated from Theorem 1).
There is a two-layer SSM that predictively models $\mathcal{L}_{FF}$ at all lengths, at finite precision.
@@ B.2 Difficulty of Representing PARITY
Theorem 12 (Restated from Theorem 2).
No SSM satisfying nonnegative can recognize PARITY at arbitrary input lengths with finite precision.
@@
If each $\alpha_{j}<1$, then $u_{t}$ converges. If some $|\alpha_{j}|\geq 1$, there may be some cancellation if $\alpha_{i}=\alpha_{j}$ for some $i\neq j$; cancellation can only lead to full erasure of the relevant terms or to a remaining term with the same exponent. In conclusion, each entry $u_{t}$ will again either converge to a finite value or diverge towards $\pm\infty$.
@@
If all dimensions of $\operatorname{Mix}_{1}(h_{t},x_{t})$ converge, then $\operatorname{Norm}(\operatorname{Mix}_{1}(h_{t},x_{t}))$ will also converge to a scaled version of $\frac{\beta_{i}}{1-\alpha_{i}}$, scaled by a bounded factor as $\beta_{i}\neq 0$. Now assume some dimensions of $\operatorname{Mix}_{1}(h_{t},x_{t})$ do not converge; in this case, for any two dimensions $i,j$, either their ratio will converge to a constant, or converge to 0 or $\pm\infty$. After applying $\operatorname{Norm}(\cdot)$, the entries asymptotically dominating the others will converge to a finite value bounded, in absolute value, by 1; the others will converge to zero.
@@
For any non-star-free regular language $\mathcal{L}$, there are words $u,v,w$ such that the membership $uv^{n}w\in\mathcal{L}$ is determined by the value of $n$ modulo some finite integer $k$ (depending on $\mathcal{L}$) [mcnaughton1971counter]. Fix any such $u,v,w\in\Sigma^{*}$.
@@
We now simply absorb the operation $X\mapsto\psi_{|w|}\circ X+\gamma_{|w|}$ into $\operatorname{Mix}_{2}$, and obtain by the same arguments as in the proof of Theorem 2 that $z_{t|v|+|w|}^{(L)}$ converges as $r\rightarrow\infty$. This is a contradiction to the claim that the value of $t$ can be read out, modulo $k$, from $z_{t|v|+|w|}^{(L)}$ at finite precision. ∎
Remark 15.
As outlined in our analysis, the assumptions in Theorem 2 are based on layer-wise operations that are either linear or based on the GLU or SwiGLU activation functions. This assumption is critical to the proof: one could design activation functions that make PARITY expressible.
Given a sequence $\mathbf{x}=x_{1},\dots,x_{T}$, consider the function $f(\mathbf{x})=\frac{e^{i\pi\sum_{i=1}^{n}x_{i}}+1}{2}$. This continuous function is designed to satisfy the condition that, for bit-strings $\mathbf{x}$, $f(\mathbf{x})=1$ if $\sum_{i=1}^{n}x_{i}$ is even, and $f(\mathbf{x})=0$ otherwise. At first glance, it seems like this function can be approximated by a cumulative sum layer in combination with a two-layer SSM to compute $f(x)=\frac{e^{i\pi x}+1}{2}$.
However, this construction cannot be implemented under the condition for which we prove Theorem 2. This is because computing this function $f(x)$ inherently requires a layer-wise nonlinear operation (such as a MLP) capable of representing sine and cosine functions over arbitrarily large input values. Importantly, achieving a construction that works for any input length requires the ability to handle arbitrarily large inputs within a single operation.
A single GLU or SwiGLU activation function, or even a more classical MLP with ReLU or sigmoid activations, is not expected to represent sine and cosine functions over unbounded inputs. The reason for this limitation lies in the universal approximation results for feedforward networks. These results generally guarantee approximation within compact convergence on bounded sets, such as in the compactification of $\mathbb{R}$ or in $L^{p}$ spaces, as described in cybenko1989approximation, ito1992approximation, and arora2016understanding. None of these results extend to uniform approximation of sine or cosine over the entire real line.
Recent work by van2024noncompact addresses the universal approximation capabilities in the space $C_{b}(\mathbb{R})$, which is the class of bounded continuous functions over $\mathbb{R}$. This result is particularly relevant since approximating sine and cosine functions uniformly over $\mathbb{R}$ would fall under this category. According to their Proposition 5.5, sine and cosine functions cannot be uniformly approximated using certain activation functions, limiting the feasibility of approximating $f(x)=\frac{e^{i\pi x}+1}{2}$ in a typical MLP architecture.
Thus, it is unrealistic to expect a typical MLP, with ReLU or sigmoid activations, to implement the function $f(x)=\frac{e^{i\pi x}+1}{2}$ uniformly for arbitrarily large inputs. Consequently, a construction based on such a function would either necessitate custom activation functions, such as periodic activations specifically designed to handle sine and cosine, or require the size of the model to scale with the input length. Either solution removes apparent contradiction with Theorem 2, as these adjustments fall outside the scope of the assumptions made in our proof.
wang2024state and orvieto2024universality provide universal approximation guarantees for SSMs, but these guarantees depend on the size of the approximating network growing with input length. This dependency is clearly stated in Proposition 3.6 and Proposition 3.9 of wang2024state, and further emphasized by orvieto2024universality. in their Remark 2. Our results, in contrast, pertain to the existence of a single SSM capable of recognizing a formal language for any input length, independent of network size. Thus, such universal approximation results do not undermine Theorem 2.
@@ B.3 Proof of Theorem 4
Our proof of Theorem 4 will rely on the algebraic theory of finite automata, specifically the cascade product and the Krohn-Rhodes Theorem [krohn1965algebraic]. These techniques, originally developed in the 1960s, have recently been introduced to the theoretical study of transformers by liu2022transformers; we provide self-contained definitions and somewhat different notation, tailored to our proofs about state-space models. In general, we will find that the properties of state-space models allow more natural and directly length-generalizing implementations of these algebraic notions than what is possible for transfomers.
Recall the definition of a finite-state-automaton (Definition 7). Our construction will build on an important operation on automata, the cascade product [krohn1965algebraic, eilenberg1974automata, ginzburg2014algebraic]:
@@
While the formal definition is cumbersome, the intuition behind it is simple: The cascade product corresponds to first reading a word $\bf w$} with $\mathcal{A}_{1}$, recording the state sequence $q_{0},q_{1},\dots,q_{|{\bf w}|}\in Q_{1}$ and – at each $t=1,\dots,|{\bf w}|$ – pasting the state $q_{t-1}$ together with the input symbol $w_{t}\in\Sigma_{1}$ – resulting in a word over a new alphabet $Q_{1}\times\Sigma_{1}$, and then running $A_{2}$ on the resulting word. The overall state of $A_{2}\wr A_{1}$ after reading a word is the tuple of the states reached by $A_{2}$ and $A_{1}$. Note that we write $A_{2}\wr A_{1}$, rather than, $A_{1}\wr A_{2}$, because the second argument of the cascade product ($A_{1}$) intuitively reads the input first, preprocessing it for the other automaton, $A_{2}$ – the cascade product can thus be viewed as a kind of function composition.
@@
Fact 17 (Consequence of Krohn-Rhodes Theorem [krohn1965algebraic] and Schützenberger’s Theorem [schutzenberger1965finite]).
@@
This result follows from the Krohn-Rhodes decomposition theorem [krohn1965algebraic], which states that any finite-state automaton can be expressed as an iterated cascade product of simple automata, specifically finite simple groups and reset automata. Moreover, Schützenberger’s Theorem [schutzenberger1965finite] characterizes star-free regular languages as those whose syntactic monoids are aperiodic, meaning they contain no nontrivial groups. Therefore, the decomposition for star-free languages involves only set-reset automata, leading to the stated cascade product structure. We now formally show that cascade products can be translated to SSM stacking. We need an auxiliary lemma, which provides a single-layer SSM that encodes the input $w_{t-1}$ in state $h_{t}$ – we will use it to forward information about the state of $\mathcal{A}_{1}$ at t1t-1 to $\mathcal{A}_{2}$ at $t$:
@@
Now we separately consider the state $h_{t}$ depending on the form of the prefix $w_{1\dots t}$ (here $w_{1\dots t}$ refers to first $t$ characters in the word). If $w_{1\dots t}=\dots 00$ (the last 2 characters of the prefix are $00$), then
@@
because
$h_{t}=$ $A(e_{0})\circ h_{t-1}+B(e_{0})$
$=$ $A(e_{0})\circ A(e_{0})\circ h_{t-2}+A(e_{0})\circ B(e_{0})+B(e_{0})$
$=$ $[1/16,1/16,0,0]^{T}\circ h_{t-2}+[1/16,1/16,0,0]^{T}\circ[1,0,1,0]^{T}+[1,0,1,0]^{T}$
$=$ $\left(\begin{matrix}\frac{1}{16}(h_{t-2})_{1}+\frac{1}{16}+1\\ \frac{1}{16}(h_{t-2})_{2}\\ 1\\ 0\end{matrix}\right)$
By definition of $A$ and $B$, each entry in $h_{t-2}$ is in $[0,2]$; the claim (14) then follows. If $w_{1\dots t}=\dots 10$, then, by a similar calculation
@@
Now, for each $\sigma\in\Sigma$, we run such a one-layer SSM where 0 represents $\sigma$ and $1$ represents all other characters. [^7] In fact, using a binary encoding of $\Sigma$, one can achieve $d=4\log|\Sigma|$. By running these in parallel (i.e. executing these operations with the same SSM layer simultaneously, utilising the width of the SSM layer) we obtain an SSM with $d=4|\Sigma|$ from whose states one can read out $w_{t-1}$ at finite precision. As the entries in $h_{t}$ are all bounded by $2$, we find $\|h_{t}\|_{2}\leq 2\sqrt{d}$ independent of $t$, and the margin is still bounded away from zero after normalization, and thus in $z_{t}$, where we can assume $\operatorname{Mix}_{1}$, $\operatorname{Mix}_{2}$ to be the identity. ∎
@@
Some SSMs include local convolutions [e.g. fu2023hungry, Gu2023Mamba] or local attention [De2024Griffin], which aggregate information from a local window of some width $\Delta>0$. These do not increase the expressive capacity beyond SSMs as we have defined in (1-2), as aggregation of local information can be simulated with a single SSM layer: Using the layer constructed in the proof of Lemma 18, given the state $h_{t}$, once one has read out $w_{t-1}$ as described in the proof, one can recover $h_{t-1}$ from $h_{t}$ and $x_{t}$; then inductively read out $w_{t-2}$ using $h_{t-1}$ and $x_{t-1}$, etc. Thus, up to any given width $\Delta>0$, one can read out $w_{t-\Delta},\dots,w_{t-1}$ from the state $h_{t}$ of this layer at finite precision.
@@
We note that a conceptually related result holds for transformers [Lemma 12 in liu2022transformers]. However, SSMs allow a simpler and length-independent construction, as they do not require positional encodings to implement such a construction.
@@
Theorem 21 (Restated from Theorem 4).
Let $\mathcal{L}$ be a regular language. The following are equivalent:
  1. 1.
    There is an SSM satisfying nonnegative that predictively models $\mathcal{L}$ at all input lengths, at finite precision
  2. 2.
    $\mathcal{L}$ is star-free.
@@
Assume $\mathcal{L}$ is star-free. By the Krohn-Rhodes theorem, there is an automaton $\mathcal{A}$ that is a cascade product of some set-reset automata that recognizes $\mathcal{L}$. By Lemmas 9 and 20, there is an SSM that computes the state sequence of that automaton.
@@
The set of regular languages known to be in $\operatorname{TC}^{0}$ is the set of regular languages whose syntactic monoid contains no non-solvable groups [barrington1992regular]. These languages are recognized by cascade products of set-reset automata and automata perfoming modular counting [straubing2012finite]. By the remark above, together with Lemma 9 and Lemma 20, such cascade products can be simulated by SSMs. ∎